# UFID: A Unified Framework for Input-level Backdoor Detection on Diffusion Models

### v1

@@ Appendix A Pseudo code of the UFID Detection Algorithm
  Input: User Input $x_{i}$; Target Diffusion Model $f_{\theta}$; Detection threshold $\tau$.
@@
     $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}+\epsilon_{j},\forall 1\leq j\leq|\mathbb{M}|\}$
@@ Appendix B More Details about Attack Baselines
VillanDiffusion [15].
We implement VillanDiffusion following the public code [^3] https://github.com/IBM/VillanDiffusion/tree/main on GitHub. As described, VillanDiffusion is a general framework for injecting backdoors into either conditional diffusion models or unconditional diffusion models. In this paper, we use VillanDiffusion specially refer to the backdoor attacks on conditional diffusion models. Specifically, the backdoor attacks is conducted over a pre-trained stable diffusion model [^4] https://huggingface.co/CompVis/stable-diffusion-v1-4/tree/main, so as to make the model generates target images once the caption trigger appears. We use "mignneko" as the caption trigger, and the Cat image used in the original paper as the target image, since these configurations are shown to perform well on different datasets in the original paper. We give an illustration of VillanDiffusion in the Figure 12.
@@
Rickrolling [42].
We implement Rickrolling following the public code [^5] https://github.com/LukasStruppek/Rickrolling-the-Artist on GitHub. As described, Rickrolling injects backdoors into the text encoder, by making the text encoder consistently generate the embedding of a target text when the trigger is present. It uses the Cyrillic o as the trigger and replaces $o$ in the original text to construct backdoor samples. The target text is chosen as "a drawing of a bird with blue eyes". We give a illustration of Rickrolling in the Figure 13.
@@
Appendix C More Details about How to Choose $\tau$.
Suppose we are given $n$ clean validation samples: $x_{1}$, $x_{2}$, …, $x_{n}$, then we take them as a batch and query the diffusion model as described in 2. In this way, a similarity graph $G$ can be constructed on this batch, with each edge denoting the similarity between any two generated images. Finally, for each node, we calculate an average similarity between this node to the other nodes. The maximal average value is used as the threshold $\tau$.
@@ Appendix D More Ablation Studies on Pre-trained Encoders
Table 4: Performance of our detection method against TrojDiff(Out-D2D) on the Cifar10 dataset with different pre-trained encoders.
@@
Table 5: Performance of our detection method against TrojDiff(D2I) on the Cifar10 dataset with different pre-trained encoders.
@@
Table 6: Performance of our detection method against BadDiffusion on the Cifar10 dataset with different pre-trained encoders.
@@ Appendix E More Details about Similarity Graphs
We provide additional qualitative examples of similarity graphs in Figure 14, Figure 15 and Figure 16. Specifically, Figure 14 presents similarity graphs for backdoor attacks on the Cifar10 dataset, where the leftmost image represents a similarity graph for a clean query. Moving from left to right, we present qualitative examples of similarity graphs for backdoor queries under TrojDiff(D2I), TrojDiff(Out-D2D), and BadDiffusion, respectively. Moreover, Figure 15 and Figure 16 present similarity graphs for VillanDiffusion and Rickrolling backdoor attacks, where the left image represents a similarity graph for a clean query, and the right image is a similarity graph for a backdoor query.
Figure 14: Similarity graphs generated for backdoor attacks on the Cifar10 dataset. The leftmost image represents a similarity graph for a clean query. Moving from left to right, we present qualitative examples of similarity graphs for backdoor queries under TrojDiff(D2I), TrojDiff(Out-D2D), and BadDiffusion, respectively.
@@
Appendix F More Details about Score Distributions
@@
Figure 17: Distributions of detection scores for backdoor samples and clean samples against unconditional diffusion models.
@@
Appendix G Proof of Theorem 3.3
Theorem G.1.
@@
Given diffusion model $f_{\theta}$, clean input noise $x_{T}\sim\mathcal{N}(0,I)$, and backdoor input noise $x_{T}^{b}=x_{T}+\delta\sim\mathcal{N}(\delta,I)$, if we perturb the clean input noise and backdoor input noise with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then the generated images from clean input noise will be more diversified than that from backdoor input noise.
@@
Definition G.2 (Clean Forward process).
Let $x_{0}\sim q(x)$ denote a sample from the clean data distribution, $x_{T}\sim\mathcal{N}(0,I)$ denote the pure Gaussian noise. Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM [22], define the forward process to diffuse $x_{0}$ to $x_{T}$ for clean samples:
@@
Definition G.3 (Clean Reverse process).
@@
Detailed proof can be found in [24]. In our setting, we assume that the attacker is able to deploy a well-trained diffusion model on the internet. Accordingly, we make the following assumptions:
Assumption G.4.
@@
This assumption implies that the noise predictors $\epsilon_{\theta}$ and $\epsilon_{\tilde{\theta}}$ are well-trained to accurately estimate the noise required to derive $x_{t}$ and $x^{\prime}_{t}$, respectively. As a result, both $f_{\theta}$ and $f_{\tilde{\theta}}$ can generate images following clean data distribution $q(x)$, given inputs following $\mathcal{N}(0,I)$ and $\mathcal{N}(0,\rho^{2}I)$, respectively. The forward and backward processes of $f_{\theta}$ are already defined from Equation 3 to 7. Notably, in our analysis, $\rho^{2}$ is set to 2 for clean samples to account for the addition of Gaussian noise. Hence, for $f_{\tilde{\theta}}$, the forward process is:
@@
$\begin{split}q(x^{\prime}_{t-1}|x^{\prime}_{t},x_{0})&=q(x^{\prime}_{t}|x^{\prime}_{t-1},x_{0})\frac{q(x^{\prime}_{t-1}|x_{0})}{q(x^{\prime}_{t}|x_{0})}\\ &\propto\exp\Big{(}-\frac{1}{2}\big{(}\frac{(x^{\prime}_{t}-\sqrt{\alpha_{t}}x^{\prime}_{t-1})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{(x^{\prime}_{t-1}-\sqrt{\bar{\alpha}_{t-1}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2}\big{(}\frac{{x^{\prime}}_{t}^{2}-2\sqrt{\alpha_{t}}x_{t}x_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\alpha_{t}}{x^{\prime}}_{t-1}^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{{x^{\prime}}_{t-1}^{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{-2\sqrt{\bar{\alpha}_{t-1}}x_{0}}x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\bar{\alpha}_{t-1}x_{0}^{2}}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}\big{(}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}){x^{\prime}}_{t-1}^{2}-(\frac{2\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{2\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+C(x^{\prime}_{t},x^{\prime}_{0})\big{)}\Big{)}}\\ &:=\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)),\end{split}$ (10)

### v2

@@ Appendix A Pseudo code of the UFID Detection Algorithm
  Input: User Input $x_{i}$; Target Diffusion Model $f_{\theta}$; Detection threshold $\tau$; small weight $\alpha=0.01$.
@@
     $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}+\alpha\cdot\epsilon_{j},\forall 1\leq j\leq|\mathbb{M}|\}$
@@ Appendix B More Details about Attack Baselines
VillanDiffusion (Chou, Chen, and Ho 2023b).
We implement VillanDiffusion following the public code [^4] https://github.com/IBM/VillanDiffusion/tree/main on GitHub. As described, VillanDiffusion is a general framework for injecting backdoors into either conditional diffusion models or unconditional diffusion models. In this paper, we use VillanDiffusion specially refer to the backdoor attacks on conditional diffusion models. Specifically, the backdoor attacks is conducted over a pre-trained stable diffusion model [^5] https://huggingface.co/CompVis/stable-diffusion-v1-4/tree/main, so as to make the model generates target images once the caption trigger appears. We use ”mignneko” as the caption trigger, and the Cat image used in the original paper as the target image, since these configurations are shown to perform well on different datasets in the original paper. We give an illustration of VillanDiffusion in the Figure 9.
@@
Rickrolling (Struppek, Hintersdorf, and Kersting 2022).
We implement Rickrolling following the public code [^6] https://github.com/LukasStruppek/Rickrolling-the-Artist on GitHub. As described, Rickrolling injects backdoors into the text encoder, by making the text encoder consistently generate the embedding of a target text when the trigger is present. It uses the Cyrillic o as the trigger and replaces $o$ in the original text to construct backdoor samples. It supports two attack modes: Target Prompt Attacks (TPA) and Target Attribute Attacks (TAA), respectively. For the TPA, the target text is chosen as ”a drawing of a bird with blue eyes”, while for the TAA, the target text is chosen as ”black and white photo”. We give a illustration of Rickrolling in the Figure 10.
@@
Personalization (Huang et al. 2024)
We implemented Personalization following the instructions [^7] https://github.com/Huang-yihao/Personalization-based__\__backdoor on Gihutb. Our implementation follows the pipeline in textual inversion, but we construct the training dataset with mismatched text-image pairs. Specifically, we use $\langle dog\rangle$ as the trigger and the cat toy images [^8] https://huggingface.co/datasets/valhalla/images as the target images. The Personalization is originally implemented to directly attack pre-trained Stable Diffusion models. To align with our experimental settings, we fine-tune the Stable Diffusion models on the Pokemon/CelebA-D dataset before conducting backdoor injection.
Appendix C More Implementation Details
Overall Implementation
All the models are well-trained with the default hyper-parameters reported in the original papers so that they show a good performance in generating both clean images and backdoor images. Following the previous works (Lee et al. 2018; Guo et al. 2023), we then evaluate our detection method with a positive (i.e., attacked) and a negative (i.e., clean) dataset. For evaluations against unconditional diffusion models, we randomly generate 1000 Gaussian noises as the clean queries (negative) and construct backdoor samples (positive) accordingly by blending the trigger pattern with the Gaussian noises. For evaluations against conditional diffusion models, we split the whole dataset into 90% train and 10% test following (Chou, Chen, and Ho 2023b). Then, we use the textual caption in the test subset as the clean queries (negative) and construct backdoor queries (positive) accordingly. Following a practical assumption in backdoor detection (Guo, Li, and Liu 2021), the threshold value $\tau$ is determined by a small clean hold-out validation dataset, where detailed descriptions are provided in the next section. The pre-trained encoder is set as CLIP-ViT-B32 (Radford et al. 2021), the size of magnitude set is chosen as 4, the poisoning rate is set as 10%, and the number of validation datasets is set as 20, by default. All the hyperparameters are evaluated in the ablation studies.
More Details about How to Choose $\tau$.
Suppose we are given $n$ clean validation samples: $x_{1}$, $x_{2}$, …, $x_{n}$, then we take them as a batch and query the diffusion model as described in 3. In this way, a similarity graph $G$ can be constructed on this batch, with each edge denoting the similarity between any two generated images. Finally, for each node, we calculate an average similarity between this node to the other nodes. The maximal average value is used as the threshold $\tau$.
Computational Resources
We conduct all the experiments on a server with 4×4$\times$4 × 80GB NVIDIA A100s.
@@ Appendix D More Ablation Studies on Pre-trained Encoders
Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
Precision 0.94 0.95 0.89 0.85 0.91 0.80 0.81 0.85 0.80
Recall 0.93 0.95 0.88 0.81 0.91 0.79 0.70 0.78 0.65
AUC 0.98 0.99 0.88 0.95 0.97 0.88 0.99 1.00 0.99
@@
Table 4: Performance of our detection method against TrojDiff(Out-D2D) on CIFAR10 dataset with different pre-trained encoders.
@@
Table 5: Performance of our detection method against TrojDiff(D2I) on CIFAR10 dataset with different pre-trained encoders.
@@
Table 6: Performance of our detection method against BadDiffusion on CIFAR10 dataset with different pre-trained encoders.
@@ Appendix E More Details about Similarity Graphs
We provide additional qualitative examples of similarity graphs in Figure 11, Figure 12 and Figure 13. Specifically, Figure 11 presents similarity graphs for backdoor attacks on CIFAR10 dataset, where the leftmost image represents a similarity graph for a clean query. Moving from left to right, we present qualitative examples of similarity graphs for backdoor queries under TrojDiff(D2I), TrojDiff(Out-D2D), and BadDiffusion, respectively. Moreover, Figure 12 and Figure 13 present similarity graphs for VillanDiffusion and Rickrolling backdoor attacks, where the left image represents a similarity graph for a clean query, and the right image is a similarity graph for a backdoor query.
Figure 11: Similarity graphs generated for backdoor attacks on CIFAR10 dataset. The leftmost image represents a similarity graph for a clean query. Moving from left to right, we present qualitative examples of similarity graphs for backdoor queries under TrojDiff(D2I), TrojDiff(Out-D2D), and BadDiffusion, respectively.
@@
Appendix F More Ablation Studies on Magnitude Set
Figure 14 presents the impact of the size of magnitude set on the performance against Personalization.
Figure 14: Performance with different magnitude size.
Appendix G More Ablation Studies on Poisoning Rates
Figure 15 presents the impact of the poisoning rate on the performance of the UFID against different types of backdoor attacks. Note that Personalization injects backdoors with only 3-5 samples, without any definitions of ”poisoning rate”.
Figure 15: Performance with different poisoning rates.
Appendix H More Ablation Studies on Available Samples
Figure 16 and Figure 17 present the impact of the available validation dataset on the performance of UFID.
Figure 16: Performance with different amounts of available samples.
Figure 17: Performance with different amounts of available samples.
Appendix I More Details about Score Distributions
@@
Appendix J Ablation Studies on Image Size
In Figure 20, we evaluate UFID against TrojDiff and BadDiffusion on an augmented CIFAR10 dataset, where the image size in CIFAR10 dataset is manually scaled to 64, 128, and 256. We could see that the image size does not have any influence on the performance, demonstrating UFID’s potential to handle high-resolution images.
Figure 18: Distributions of detection scores for backdoor and clean samples on unconditional diffusion models.
@@
Figure 20: Performance with different Image Size.
Appendix K Adaptive Attacks
Figure 21 shows the performance of UFID against adaptive attacks.
Figure 21: Evaluation of UFID against adaptive attacks.
Appendix L Evaluations with Diversity-intensive Backdoor Attacks.
BadT2I (Zhai et al. 2023) is a novel diversity-intensive backdoor attack method. BadT2I contains three modes: Pixel-backdoor, Object-backdoor, and Style-backdoor, respectively. For simplicity, we start our analysis based on the most difficult one, i.e., object-backdoor here. The analysis can be easily applied to the other two modes. Detailed descriptions of the other two modes can be found in the original paper.
[…]
The proposed UFID pipeline is not able to effectively distinguish backdoor generations from clean generations since the backdoor generations (i.e., bike images) can be as diversified as the clean generations. For example, in Figure 22, the graph density scores for the clean generations and backdoor generations are hardly distinguished. Despite the great challenge, we found that additional correspondence information between the input prompt and the generations could be integrated into the existing UFID pipeline without violating the black-box assumptions in the threat model.
[…]
$\texttt{Corre}(x_{i},y_{i})=-\langle\texttt{CLIP}(y_{i}),\texttt{CLIP}(x_{i})\rangle$ (4)
Figure 22: Graph density score distributions of backdoor samples and clean samples.
Figure 23: Detection score distributions of backdoor samples and clean samples.
Then the final detection score can be the sum of the Corre score and the graph density score. It is noted that the graph density score here is manually scaled with a weight $(|\mathbb{M}|-1)$ to match the scale of the Corre score, where $|\mathbb{M}|$ is the size of the generated batch.
We conduct experiments on three types of BadT2I: pixel-backdoor, object-backdoor, and style-backdoor. For each type of BadT2I, we ask ChatGPT to generate 100 random prompts according to their default specifications. The backdoor samples are constructed by concatenating the trigger with each prompt. The following Table 9 shows the AUC values of our detection performance.
Pixel-backdoor Object-backdoor Style-backdoor
UFID (w/ Corre) 0.90 0.94 0.86
Table 9: AUC values of our enhanced UFID method on three types of BadT2I backdoor attacks.
Appendix M Social Impact Statment
Appendix N Experiments about Model-free Similarity Metric
To relax the assumption of using a pre-trained encoder for calculating image similarities, we explore model-free metrics like SSIM for image similarity. Table 10 reports additional experiments on CIFAR10 with TrojDiff, demonstrating that UFID is effective with SSIM as the similarity metric.
Table 10: Effectiveness of UFID on Cifar10 dataset with SSIM.
Attack P R AUC
TrojDiff(D2I) 0.95 0.96 1.00
TrojDiff(Out) 0.93 0.91 0.98
TrojDiff(In) 0.73 0.76 0.84
Figure 24: Comparison of the problem settings.
Appendix O Proof of Lemma 1
Lemma 7.
Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in the Assumption 11. Let input noise $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$. Let $\hat{x_{0}}$ be the generated image for $x^{\prime}_{T}$.Let the clean data distribution be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$. We then have:
Proof.
the Equation 6 is due to Assumption 11. In particular, to obtain the generated image follows $q(x)$, the reverse process is defined as $q(x^{\prime}_{t-1}|x^{\prime}_{t})\sim\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t))$, where $\mu_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}{\epsilon}_{t}\Big{)}$ and $\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}$ (Equation 8). For the $f_{\theta}$, although the reverse process is also a gaussian distribution, $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(x_{t},t)\Big{)}$, $\Sigma_{\theta}(x_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$ (Equation 13 and 14).
Under the Assumption 11, $\epsilon_{\theta}$ is able to accuaratly predict the noise added on the $\sqrt{\bar{\alpha}_{t}x_{0}}$ to obtain $x_{t}$, hence the prediction of $\epsilon_{\theta}$ in Equation 7 should be $\rho\epsilon_{t}$. We have by substituting $\rho\epsilon_{t}$ into Equation 7:
By comparing Equation 9 and Equation 8 with Equation 8 , we found the mean of the reverse process is the same when inputting the $x^{\prime}_{T}$ to the $f_{\theta}$ and $f_{\tilde{\theta}}$, while the variance of $f_{\tilde{\theta}}$ is $\rho^{2}$ times larger than $f_{\theta}$. For simplicity, Let $a_{t}=\mu_{\theta}(x^{\prime}_{t},t)$ and $b_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$, we have: $q_{\theta}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}I)$ and $q_{\tilde{{\theta}}}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}I)$. Hence, by the reparameterization trick, the variance of the generated $\hat{\tilde{x_{0}}}$ of $f_{\tilde{\theta}}$ is $\rho^{2}$ times greater than $f_{\theta}$. Without loss of generality, we use the Gaussian distribution to discribe the output distribution. Given the Assumption 11, and $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$, $\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$, which completes the proof. ∎
Appendix P Proof of Theorem 2
Theorem 8.
@@
Suppose the output domain of the diffusion model $f_{\theta}$ is Gaussian. Let $\mathcal{N}(\mu_{c},\sigma_{c})$ and $\mathcal{N}(\mu_{b},\sigma_{b})$ denote the distribution of clean generations and backdoor generations respectively. Let $N$ denote the image size. We assume that $\sigma_{c}\geq\sigma_{b}+\rho^{2}$. Given clean input noise $x_{T}^{c}\sim\mathcal{N}(0,I)$, backdoor input noise $x_{T}^{b}=x_{T}^{c}+\delta\sim\mathcal{N}(\delta,I)$, and Lemma 1 if we perturb the clean input noises $x_{T}$ and backdoor input noise $x_{T}^{b}$ with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then for the resulted clean generations $\mathcal{N}(\mu^{\prime}_{c},\sigma^{\prime}_{c})$ and the backdoor generations $\mathcal{N}(\mu^{\prime}_{b},\sigma^{\prime}_{b})$, we have that $\sigma^{\prime}_{c}-\sigma^{\prime}_{b}\geq 1$.
@@
Definition 9 (Clean Forward process).
Let $x_{0}\sim q(x)$ denote a sample from the clean data distribution, $x_{T}\sim\mathcal{N}(0,I)$ denote the pure Gaussian noise. Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM (Ho, Jain, and Abbeel 2020), define the forward process to diffuse $x_{0}$ to $x_{T}$ for clean samples:
@@
Definition 10 (Clean Reverse process).
@@
Detailed proof can be found in (Hu et al. 2023). In our setting, we assume that the attacker is able to deploy a well-trained diffusion model on the internet. Accordingly, we make the following assumptions:
Assumption 11.
@@
This assumption implies that the noise predictors $\epsilon_{\theta}$ and $\epsilon_{\tilde{\theta}}$ are well-trained to accurately estimate the noise required to derive $x_{t}$ and $x^{\prime}_{t}$, respectively. As a result, both $f_{\theta}$ and $f_{\tilde{\theta}}$ can generate images following clean data distribution $q(x)$, given inputs following $\mathcal{N}(0,I)$ and $\mathcal{N}(0,\rho^{2}I)$, respectively. The forward and backward processes of $f_{\theta}$ are already defined from Equation 10 to 14. Notably, in our analysis, $\rho^{2}$ is set to 2 for clean samples to account for the addition of Gaussian noise. Hence, for $f_{\tilde{\theta}}$, the forward process is:
@@
$\begin{split}q(x^{\prime}_{t-1}|x^{\prime}_{t},x_{0})&=q(x^{\prime}_{t}|x^{\prime}_{t-1},x_{0})\frac{q(x^{\prime}_{t-1}|x_{0})}{q(x^{\prime}_{t}|x_{0})}\\ &\propto\exp\Big{(}-\frac{1}{2}\big{(}\frac{(x^{\prime}_{t}-\sqrt{\alpha_{t}}x^{\prime}_{t-1})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{(x^{\prime}_{t-1}-\sqrt{\bar{\alpha}_{t-1}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2}\big{(}\frac{{x^{\prime}}_{t}^{2}-2\sqrt{\alpha_{t}}x_{t}x_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\alpha_{t}}{x^{\prime}}_{t-1}^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{{x^{\prime}}_{t-1}^{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{-2\sqrt{\bar{\alpha}_{t-1}}x_{0}}x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\bar{\alpha}_{t-1}x_{0}^{2}}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}\big{(}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}){x^{\prime}}_{t-1}^{2}-(\frac{2\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{2\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+C(x^{\prime}_{t},x^{\prime}_{0})\big{)}\Big{)}}\\ &:=\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)),\end{split}$